2/3x - 3/4 = 1/6x + 21/4
Answer by stanbon(75887)

(2/3)x - 3/4 = (1/6)x + 21/4
Multiply thru by 4 to get:
4(2/3) - 3x = 4(1/6) + 21
(8/3) - 3x = (2/3) + 21
3-21 = 3x
-18 = 3x
x = -6
=================
Cheers,
Stan H.